Liquidating Partnerships Prior to liquidating their partnership, Todd and Dunn had capital accounts of $62,000 and $119,000, respectively. Prior to liquidation, the partnership had no cash assets other than what was realized from the sale of assets. These partnership assets were sold for $216,000. The partnership had $7,000 of liabilities. Todd and Dunn share income and losses equally.
Answer #1

Todd Dunn Total Capital Aelel: inip liabiliiy Sub Total Less: Asset Sold Profit Total Division 62,000 1,19,000 1,81,000 7,000 1,88,000 2,16,000 28,000 4,00014,000 76,000 1,33,000

As profit and loss of Todd and Dunn share equally the profit of $28000.

$28,000 * 50% = $14,000

Todd share of capital is 62,000 + 14000 = $76,000

Dunn Share Capital is 119,000 + 14,000 =  $133,000.